OpenStudy (anonymous):

using logarithmic differentiation find the derivative of the function:

OpenStudy (anonymous):

\[\sqrt[6]{(x^2 +1)/(x^2 -1)}\]

OpenStudy (anonymous):

You need a y=

OpenStudy (anonymous):

yeah y = that

OpenStudy (anonymous):

ok introduce ln to y and your function.

OpenStudy (anonymous):

i have \[\ln y = 1/6 \ln (x^2+1)/(x^2-1)\]

OpenStudy (anonymous):

Ok now remember, \[\ln(a/b) = ln a - ln b\]

OpenStudy (anonymous):

so it's \[\ln y = 1/6 \ln(x^2+1) - \ln(x^2-1) \] ?

OpenStudy (anonymous):

The 1/6 should be multiplied to both functions but we can factor it to avoid any mistakes.

OpenStudy (anonymous):

\[lny= 1/6( \ln(x^2+1) - \ln (x^2-1))\]

OpenStudy (anonymous):

Now can you go on?

OpenStudy (anonymous):

no note quite.....

OpenStudy (anonymous):

ok,\[ d/dx lnx= 1/g(x) \times g'(x)\] right?

OpenStudy (anonymous):

\[d/dx (lnx)=1/g(x)×g′(x)\]

OpenStudy (anonymous):

yes d/dx [ln g(x)] = g'(x)/g(x)

OpenStudy (anonymous):

ok so differentiate each term using that rule first and remember the derivative of y is y' since we are looking for dy/dx. Then let me know what you get and we can move on to the next step.

OpenStudy (anonymous):

so would it be \[\ln y =2x/[(x^2+1)/(x^2-1)] d/dx[(x^2+1)/(x^2-1)] \] all times 1/6 ?

OpenStudy (anonymous):

I'm not sure why you have a d/dx there. But this should be simple. Let me do half the question for you. derivative of\[ lny = 1/y \times dy/dx\] Derivative of\[ \ln (x^2+1) = 1/(x^2+1) \times 2x\] Derivative of\[ \ln (x^2-1) = 1/(x^2-1) \times 2x\] Go on from here and let me know what you have as dy/dx

OpenStudy (anonymous):

\[d/dx[\ln(x^2+1)-\ln(x^2-1)]1/6 = 1/6[(2x/2x+1)-(2x/x^2-1)]\]

OpenStudy (anonymous):

Hold on, It's \[x^2+1\] not \[2x+1\]. Also, the question is not yet finished because \[1/y \times dy/dx= (2x/x^2+1)- (2x/x^2-1)\] So we are meant to solve for dy/dx. What can you do to solve for dy/dx?

OpenStudy (anonymous):

i can use chain rule to make \[dy/dx = dy/du *du/dx = 1/u du/dx\]

OpenStudy (anonymous):

No you do not have to, since we are looking for dy/dx. we just multiply everything by y to get \[dy/dx=1/6y(2x/x2+1)−(2x/x2−1)\] But still we are not done.

OpenStudy (anonymous):

i'm sorry i'm still completely clueless

OpenStudy (anonymous):

ok, when they ask you to differentiate using logs, they want you to find dy/dx

OpenStudy (anonymous):

So what you do is IMPLICITLY differentiate everything.

OpenStudy (anonymous):

remember implicit differentiation?

OpenStudy (anonymous):

i'm trying to. my book isn't being of much help

OpenStudy (anonymous):

You see y' is the same as dy/dx

OpenStudy (anonymous):

In the question I just typed, to find y' or dy/dx. We would have to differentiate all terms.

OpenStudy (anonymous):

Find y' if \[2=2x^3y^2+xy\] <-- this was my question. Let's start with the left hand side, derivative of 2 is 0 Then we move on to the right hand side

OpenStudy (anonymous):

ok so that's dydx 2= 2x^3 * y^2 +xy is 0 = 6x^2*2y+1

OpenStudy (anonymous):

we have to functions that need to be derived using product rule.

OpenStudy (anonymous):

two*

OpenStudy (anonymous):

That's incorrect. Remember product rule= \[f'g + fg'\]

OpenStudy (anonymous):

\[0 = 6x^2y^2 + 2x^3*2y + y+x\]

OpenStudy (anonymous):

Ok now this is almost correct. Just remember, every time you are looking for y' or dy/dx when you differentiate y you get y' or dy/dx So the derivative of y^2= 2yy' and derivative of y= y'

OpenStudy (anonymous):

Whenever you differentiate y add a y' or dy/dx

OpenStudy (anonymous):

ok so it's \[0=6x^2y^2 + 2x^3*2yy' +y'+x\]

OpenStudy (anonymous):

and we are to isolate y'

OpenStudy (anonymous):

exactly and then solve for y'.

OpenStudy (anonymous):

But then it would be \[0=6x^2y^2+2x^3(2yy)+y+xy'\]

OpenStudy (anonymous):

sorry that's 2yy'

OpenStudy (anonymous):

\[(-6x^2y^2 - x )/ (2x^3*2y*2) = y'\]

OpenStudy (anonymous):

ok then with that infon then it's different from that

OpenStudy (anonymous):

I don''t think you isolated y' correctly. 6x^2y^2 and y would go to the left hand side of the equation so that you are left with terms that contain y' only. Then just factor out y' Try it again.

OpenStudy (anonymous):

i still get \[(-6x^2y^2 -y)/(2x^3+x+2y+2) = y'\]

OpenStudy (anonymous):

You are almost write. Are you isolating the equation I wrote down or yours? Because you were missing and xy' and y. Compare the two equations. But anyway, seems like you got the point. Now we apply the same knowledge to the log function.

OpenStudy (anonymous):

the one you wrote....yeah i think i get it mostly now. how do i go about applying it to the log eqn?

OpenStudy (anonymous):

Ok so you got everything until the part where we simplified the log right? Because before you do anything to do with log differentiation, always simplify the function.

OpenStudy (anonymous):

yes i have \[1/y*dy/dx=(2x/x^2 +1)−(2x/x^2 −1) \]

OpenStudy (anonymous):

Great, don't forget the 1/6

OpenStudy (anonymous):

So now just solve for dy/dx as we did in the previous example.

OpenStudy (anonymous):

ok so i muliply the function by y to give [dy/dx=1/6y(2x/x2+1)−(2x/x2−1)\] right?

OpenStudy (anonymous):

yes but then you would have to plug in your original function because remember. \[y= \sqrt[6]{(x2+1)/(x2−1)}\]

OpenStudy (anonymous):

so it would be \[1/6(\sqrt[6]{(x^2+1/x^2-1}) (2x/x^2+1)-(2x/x^2-1)\] ?

OpenStudy (anonymous):

It's 1/6y multiplied to the whole function so multiply \[1/6(\sqrt[6]{(x2+1/x2−1)} )\] to all of this ---->\[(2x/x^2+1)−(2x/x^2−1)\] to get \[1/6(\sqrt[6]{(x2+1/x2−1)} (2x/x^2+1)−(2x/x^2−1))\]

OpenStudy (anonymous):

that looks like what i put, is it different?

OpenStudy (anonymous):

Because \[dy/dx= 1/6y'(2x/x^2+1)−1/6y'(2x/x^2−1)\] So we just factor out the 1/6 y'

OpenStudy (anonymous):

\[1/6 \sqrt[6]{1/6((x2+1/x2−1)} ((2x/x2+1)−(2x/x2−1))\] This is your answer, notice I have double parenthesis to show that my 1/6y' is multiplied to both functions in parenthesis?

OpenStudy (anonymous):

so it is just a bit different from yours, but you get the point I see.

OpenStudy (anonymous):

I meant that is the correct answer^^ or rather the answer you need.

OpenStudy (anonymous):

so that one is the final answer or one that still needs to be simplified?

OpenStudy (anonymous):

That is your final answer, simplifying would just be tedious.

OpenStudy (anonymous):

\[ y' =( 24x/x^2 - 1 ) \sqrt[6]{(x^2+1)/(x^2-1)}\] would that translate into

OpenStudy (anonymous):

i think it might giving the factoring of the 6y' and the two 2x's

OpenStudy (anonymous):

The two 2x's have a different base so you would have to find a common denominator in order to add them.

OpenStudy (anonymous):

Also, you were missing 1/6 in that answer.

OpenStudy (anonymous):

i meant x^4 - 1 in that first part

OpenStudy (anonymous):

Your common denominator is write, but adding the 2x's would give you -4x wouldn't it? divided by \[6(x^4-1)\] because of the 1/6. So you would end up with -\[2x/3x^4-3\]

OpenStudy (anonymous):

right*

OpenStudy (anonymous):

\[-2x/3x^4-3\]

OpenStudy (anonymous):

you are right! thank you so much! you help al lot

OpenStudy (anonymous):

No problem. Just do lots of practice and you will be used to this in no time.

OpenStudy (anonymous):

:) you deserve all your medals

OpenStudy (anonymous):

haha ,you too, for your patience and interest in learning.